OpenStudy (anonymous):

Help Identify the vertex and the axis of symmetry of the graph function y=3(x+2)^2-3. V:2,-3 Axis: x=2 V:-2,-3 Axis: x=-2 V:2,3 Axis: x=2 V:-2,3 Axis:x=2

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

So, you have a parabola already in its vertex form: \(y=a(x-h)^2+k\) And the vertex of the parabola is the coordinates \((h, k)\) , can you identify h and k here?

OpenStudy (anonymous):

h=2, k=3?

OpenStudy (anonymous):

or does k=2-3

OpenStudy (kirbykirby):

Let's see if we can re-write the given equation so that it matches the format of the vertex form: \(y=3(x-(-2))^2+(-3)\), this is equivalent to the equation you were given

OpenStudy (anonymous):

Mhmm

OpenStudy (anonymous):

So so, it's V:-2,-3 Axis: x=-2?

OpenStudy (kirbykirby):

yes absolutely :)

OpenStudy (anonymous):

Thanks so much. <3 ~~[*~*]~~

OpenStudy (kirbykirby):

your welcome

OpenStudy (anonymous):

That was my first guess too! Dun worry it was an educated guess.

OpenStudy (kirbykirby):

=]